# Towards Principled Assessment of Tabular Data Synthesis Algorithms

### v1

@@ 2.2 Evaluation Pipelines
We propose SynMeter, a systematic evaluation framework for synthesizers. SynMeter consists of four phases: data preparation, model tuning, model training, and model evaluation.
@@
The data preparation phase explores and preprocesses data for learning algorithms [^3] Here we assume that there are no missing values in original data. The missing values problem has been extensively studied [51], which is independent of data synthesis.. In this phase, statistical methods [69, 70, 39] select low-dimensional marginals to serve as a compact representation for capturing data dependencies. Deep generative models [65, 26, 68, 71] involve standard data processing techniques like data encoding and normalization.
@@
The goal of model tuning phase is to select optimal hyperparameters for generative models. A subset of training dataset $\mathcal{D}_{\text{train}}$ is sampled as validation data $\mathcal{D}_{\text{val}}$. The generative model $\mathsf{G}$ is trained on $\mathcal{D}_{\text{train}}$ and assessed its performance on $\mathcal{D}_{\text{val}}$ for model tuning. Note that the tuning phase is often overlooked in many data synthesis models. However, we show in Section 7.6 that a proper tuning objective can boost the performance of synthesizers.
@@
In the model evaluation phase, the trained model generates some synthetic data $\mathcal{S}$, which are used for evaluation.
3 Fidelity Evaluation
@@
In this section, we first review existing fidelity measurements and identify their limitations, and then introduce a new and more principled fidelity metric.
3.1 Existing Metrics and Limitations
@@
Low-order Statistics. Low-dimensional marginals are crucial to the quality of synthetic data and many well-established statistics are used for one (two)-way marginals evaluation:
@@
Distribution Measurements. Total Variation Distance (TVD) [70, 39, 32] and Kolmogorov-Smirnov Test (KST) [68] are used to measure the univariate distribution similarity for categorical and numerical attributes, respectively. TVD computes the absolute probabilities difference of categorical distributions, while KST examines the maximum discrepancy between cumulative distribution functions for numerical attributes. The main problem of these methods is the lack of versatility. Each type of marginal requires a distinct statistical measure, which complicates the ability to perform a principled comparison across various data types.
Correlation Statistics. Studies [12, 32, 71, 58] often utilize correlation differences for pairwise similarity evaluations. Specifically, the correlation score of any two pairs of attributes is calculated on both synthetic and real data, and the overall mean difference of correlation scores is used as fidelity. Popular correlation statistics like Theil’s uncertainty coefficient [^4] https://en.wikipedia.org/wiki/Uncertainty_coefficient, Pearson correlation, and the correlation ratio [^5] https://en.wikipedia.org/wiki/Correlation_ratio are applied for various marginal types (categorical, continuous, and mixed). However, correlation coefficients are scale-invariant, meaning they do not take into account differences in the scale (variance) or location (mean) of distributions, which would result in a misleading perception of similarity between the synthetic and real datasets.
@@
Likelihood Fitness. This metric [65] aims to directly compute the likelihood of synthetic data as fidelity by using simulated data as input datasets. Since the input data is generated from some prior probabilistic models (e.g., Bayesian networks), the likelihood of synthetic data can be derived by fitting them to the priors. However, this approach can only be applied by simulation, which limits the application for real-world datasets.
Table 2: Comparison of existing fidelity metrics. “ ✓ ” indicates the metric meets the corresponding property, and “ ✗ ” denotes otherwise.
Fidelity Metric Type Faithfulness Universality
TVD/KST Statistics
Correlations Statistics
Likelihood Fitness Simulation
$\alpha$-Precision Coverage-based
Probabilistic MSE Model-based
Wasserstein Distance (ours) Statistics
@@
Other Approaches. $\alpha$-Precision [2] defines fidelity as the proportion that the generated sample is covered by real data. The coverage is determined by comparing the distance between synthetic data and the nearest real data to a target real one. Probabilistic mean squared error (pMSE) [56] employs a logistic regression discriminator to distinguish between synthetic and real data, using relative prediction confidence as a fidelity metric. However, these heuristic approaches are unable to accurately capture the distribution difference between synthetic data and real data.
Discussion. We compare existing metrics against the principles of fidelity evaluation, as outlined in Table 2. The results indicate that none of these metrics fulfill all the desired properties. This gap underlines the necessity for developing a new fidelity metric for holistic assessment.
3.2 Proposed Metric: Wasserstein Distance
@@
We opt for Wasserstein distance to measure the distribution discrepancies between synthetic data and real data. Originating from optimal transport theory [50], the Wasserstein distance provides a structure-aware measure of the minimal amount of work required to transform one distribution into another. Formally, Let $\mathbf{P}=(p_{1},p_{2},\ldots p_{n})$ and $\mathbf{Q}=(q_{1},q_{2},\ldots q_{n})$ be the two probability distributions, and $\mathbf{C}$ be a matrix of size $n\times n$ in which $\mathbf{C}_{ij}\geq 0$ is the cost of moving an element $i$ of $\mathbf{P}$ to the element $j$ of $\mathbf{Q}$ ($\mathbf{C}_{ii}=0$ for all element $i$). The optimal transport plan $\mathbf{A}$ is defined as the minimal amount of work between two distributions:
@@
$\begin{split}&\min_{\mathbf{A}}\quad\langle\mathbf{C},\mathbf{A}\rangle\\ \text{ s.t. }&\mathbf{A}\mathds{1}=\mathbf{P},\quad\mathbf{A}^{\top}\mathds{1}=\mathbf{Q}\end{split}$ (1)
@@
$\mathcal{W}(\mathbf{P},\mathbf{Q})=\langle\mathbf{C},\mathbf{A}^{*}\rangle$ (2)
@@
Definition 1
(Wasserstein-based Fidelity Metric) Let $\mathbf{P}$ and $\mathbf{Q}$ be the marginal distributions derived from synthetic data $\mathcal{S}$ and real data $\mathcal{D}$, the fidelity of data synthesis algorithm $\mathsf{A}$ is:
$\mathrm{Fidelity}(\mathsf{A})\coloneqq\underset{\begin{subarray}{c}\mathbf{P}\sim\mathcal{S}\\ \mathbf{Q}\sim\mathcal{D}\end{subarray}}{\mathbb{E}}\big{[}\mathcal{W}(\mathbf{P},\mathbf{Q})\big{]}$ (3)
@@
In practice, the dataset $\mathcal{D}$ can be designated as either $\mathcal{D}_{train}$ or $\mathcal{D}_{test}$, depending on the specific requirements of the analysis. We compute the Wasserstein distance of all the one-way and two-way marginals and use the mean as the final fidelity score.
Determining Cost Matrix. The Wasserstein distance requires the predefined cost matrix $\mathbf{C}$, which encapsulates the “difficulty” or “cost” of transitioning from one distribution element to another. For $k$-way marginal distributions $\mathbf{P}$ and $\mathbf{Q}$, the cost matrix is formulated by summing the pairwise distances between corresponding elements:
$\mathbf{C}_{ij}=\sum\nolimits_{r}^{k}d(v_{i}^{r},v_{j}^{r}).$ (4)
Here, $v_{i},v_{j}\in\mathbb{R}^{k}$ are the element located in $i$ and $j$ in $k$-way probability distributions, respectively. The distance $d(\cdot,\cdot)$ is tailored to the nature of the elements, differing for numerical and categorical values:
$d(v_{i}^{r},v_{j}^{r})=\begin{cases}||v_{i}^{r}-v_{j}^{r}||_{1}&\text{if numerical}\\ \infty,&\forall v_{i}^{r}\neq v_{j}^{r},\text{if categorical}.\end{cases}$ (5)
This means we use $l_{1}$ distance for numerical values and consider the strict match for categorical values (i.e., cost for any different categorical values is infinity).
Solving Wasserstein Distance. The computation of Wasserstein distance involves solving the optimization problem in Equation 1, which is an instance of linear programming problems (LP) [50]. There are many open-source libraries like CVXPY [14] and OPT [18] that can be used to solve LP reasonably fast. Besides, it can also be computed by solving an instance of minimum cost flow problem using the network simplex algorithm, which is much faster than using the LP solver when the cost matrix $\mathbf{C}$ is sparse [36].
However, when the cost matrix is large and dense, computing this metric can be prohibitively time-consuming. Several options are provided to address this problem. (i) Sinkhorn distance [13] provides a fast approximation to the Wasserstein distance by penalizing the objective with an entropy term. (ii) Sliced-Wasserstein distance [6], which uses Radon transform to linearly project data into one dimension, can be efficiently computed. (iii) Reducing the size of the cost matrix by randomly sampling a small set of points from the probability densities. These computational strategies ensure the Wasserstein distance remains a feasible solution for fidelity evaluation, even for complex and high-dimensional marginals.
Advantages of Wasserstein-based Fidelity Metric. Wasserstein distance offers several distinct advantages for fidelity evaluations: (i) It serves as a natural and structure-aware measurement for analyzing distribution discrepancies, which generalizes existing metrics such as TVD and contingency similarity [49, 68]. (ii) It can accommodate both numerical and categorical attributes under the same criterion, which facilitates the evaluation of heterogeneous types of marginals.
4 Privacy Evaluation
@@
4.1 Existing Metrics and Limitations
@@
Similarity-based Privacy Metrics. Similarity-based metrics have become the conventional privacy evaluation metrics for HP synthesizers [38, 19, 29]. These heuristic methods empirically assess the privacy risks by measuring the closeness between synthetic data and real data. Among all heuristics, the Distance to Closest Records (DCR) is the most widely used both in academia [71, 66, 32, 68] and industry [4, 23, 42]. DCR looks at the distribution of the distances from each synthetic data point to its nearest real one, and uses the 5th percentile (or the mean) of this distribution as the privacy score. A small score indicates that the synthetic dataset is too close (similar) to real data, signaling a high risk of information leakage.
Although the prevalence of these metrics as privacy evaluation, the meaningfulness of these heuristics has never been rigorously studied. The commonly accepted concept of privacy breach is the ability of an adversary to either re-identify or assert the membership of an individual in real data [35, 16, 61, 45]. Hence, a proper privacy metric should protect every training instance against membership disclosure. However, this cannot be achieved by assessing the similarity between synthetic data and real data. To elucidate this point, consider a simple synthesizer that generates data by randomly adding a constant to the real attributes or aggregating some attributes’ values. In such a case, the similarity-based metrics may misleadingly indicate a low privacy risk since the synthetic data does not replicate any individual record. However, the synthesizer remains vulnerable since the membership disclosure risk of the aggregated records is alarmingly high.
@@
Another concern of the similarity-based privacy metrics is the risk of likability. Since these metrics are computed by comparing the closeness with real training data, using these metrics in the synthesizing process (e.g., trying to decrease DCR during synthesis) provides another channel to leak information about the training data. For example, recent research [19] has demonstrated the feasibility of reconstructing significant portions of training data, when given black-box access to a data synthesizer and an oracle for computing the DCR of synthesized datasets. This revelation serves as a stark warning to practitioners who rely on such metrics.
@@
Membership Inference Attacks. Another way to empirically verify the privacy of models is membership inference attacks [55], which has been extensively studied on discriminative models. However, membership inference is generally harder to mount on generative models than it is on discriminative ones since there are no clear confidence signals to detect overfitting and infer membership [25, 62]. Some studies [57] try to facilitate the problems by utilizing handcrafted features extracted from synthetic data distribution to train the shadow models and attack models. Another work [28] approximates the likelihood of target records with density estimation and uses it as the confidence score for attacks. However, as pointed out in [21] and [53], these attacks either depend on unrealistic settings or are too weak to differentiate the membership disclosure risks among various synthesizers [15]. Therefore, existing synthesis algorithms rarely use membership inference attacks for privacy evaluation [53].
4.2 Proposed Metric: Membership Disclosure Score
We propose a new privacy metric for data synthesis, termed membership disclosure score (MDS), to directly estimate the membership disclosure risks without relying on the effectiveness of attacks. We first quantify the disclosure risk of the target record $x$ as follows:
@@ 4.2 Proposed Metric: Membership Disclosure Score
$\mathrm{DS}(x,\mathsf{A},\mathcal{D})\coloneqq\underset{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime}\subset\mathcal{D}\\ {x\in\mathcal{H},x\notin\mathcal{H}^{\prime}}\end{subarray}}{\mathbb{E}}\big{[}||\text{NN}(\mathcal{S},x)-\text{NN}(\mathcal{S}^{\prime},x)||_{1}\big{]},$ (6)
@@
where $\mathcal{S}$ and $\mathcal{S}^{\prime}$ represent the synthetic datasets when the synthesizer $\mathsf{A}$ uses the subsets $\mathcal{H}$ and $\mathcal{H}^{\prime}$ respectively as the input to output a data generator. The term $\text{NN}(\mathcal{S},x)$ denotes the nearest neighbor (under $l_{1}$ distance) for record $x$ within the synthetic dataset $\mathcal{S}$. We also use $l_{1}$ distance [^6] Empirically we find that the difference between using $l_{1}$ and $l_{2}$ distance for computation is negligible. to measure the discrepancy resulting from training the model with or without the record $x$. Note that the disclosure score is calculated based on the difference between neighboring synthetic data, which ensures that real data information remains undisclosed. Accordingly, the membership disclosure risk of synthesizer $\mathsf{A}$ can be defined as:
Definition 2
@@
(Membership Disclosure Score) Let $\mathrm{DS}(x,\mathsf{A},\mathcal{D})$ be the disclosure risk of record $x$ in real data $\mathcal{D}$ for synthesis algorithm $\mathsf{A}$. The membership disclosure score of the data synthesis algorithm $\mathsf{A}$ is given by:
$\mathrm{MDS}(\mathsf{A})\coloneqq\max_{x\in\mathcal{D}}~{}\mathrm{DS}(x,\mathsf{A},\mathcal{D}).$ (7)
@@
Here we omit the notion of dataset $\mathcal{D}$ for simplicity. This metric signifies the maximum disclosure risk across the entire training dataset, which indicates that MDS is a worst-case privacy measurement. Intuitively, for a privacy-preserving synthesizer, the change of output (synthetic data) should remain insignificant by any individual training samples, resulting in a low MDS.
@@
Directly computing this score entails retraining the model for each individual record, which is computationally prohibitive. Alternatively, we train $m$ models once on random subsets of $\mathcal{D}$, and compute the disclosure score for all instances using the same set of $m$ synthesis models. This approach provides an efficient and practical approach to estimating MDS in real-world scenarios.
Our definition of MDS aligns closely with concepts of memorization in neural networks [17, 67] and the leave-one-out notion of stability in machine learning [8]. However, it diverges in three crucial ways: (i) We measure the worst-case discourse risk as the privacy metric, whereas other studies focus on the difference of individuals or average cases. (ii) Our work specifically addresses privacy concerns in data synthesis, as opposed to other studies that explore discriminative models like classification. (iii) Our approach emphasizes the discrepancy caused by the presence or absence of target data in training, in contrast to other works that highlight performance gains from adding samples to the training set.
5 Utility Evaluation

### v2

@@ 2.2. Evaluation Pipelines
We propose $\mathsf{SynMeter}$, a systematic evaluation framework for synthesizers. The $\mathsf{SynMeter}$ pipeline consists of four phases: data preparation, model tuning, model training, and model evaluation.
@@
The data preparation phase preprocesses data for learning algorithms. (Here we assume no missing values in the original data. The missing values problem has been extensively studied (Pigott, 2001), which is orthogonal to data synthesis.) In this phase, statistical methods (Zhang et al., 2017, 2021; McKenna et al., 2021) select low-dimensional marginals to serve as compact representations for capturing data distributions. Deep generative models (Xu et al., 2019; Ho et al., 2020; Zhang et al., 2023b; Zhao et al., 2021) apply standard data processing techniques like data encoding and normalization.
@@
The goal of model tuning phase is to select the optimal hyperparameters for data synthesizers. While this phase is often neglected in many data synthesis models, we show in Section 7.5 that a proper tuning objective can boost the performance of synthesizers.
@@
In the model evaluation phase, the trained model generates some synthetic data, which are used for evaluation.
3. Fidelity Evaluation
@@
In this section, we first review existing fidelity measurements, identify their limitations, and introduce a new fidelity metric that addresses these limitations.
3.1. Existing Metrics and Limitations
@@
Low-order Statistics. Marginals are the workhorses of statistical data analysis and well-established statistics for one(two)-way marginals have been used to assess the quality of synthetic data.
@@
Distribution Measurements. Total Variation Distance (TVD) (Zhang et al., 2021; McKenna et al., 2021; Kotelnikov et al., 2023) and Kolmogorov-Smirnov Test (KST) (Zhang et al., 2023b) are used to measure the univariate distribution similarity for categorical and numerical attributes, respectively. The main problem with this approach is the lack of versatility. Each type of marginal requires a distinct statistical measure, which complicates the ability to perform a comprehensive comparison across various attribute types.
Correlation Statistics. Some researchers use correlation difference, i.e., the difference of correlation scores on synthetic and real data, to measure the pairwise distribution similarity. Popular correlation statistics like Theil’s uncertainty coefficient (Zhao et al., 2021), Pearson correlation (Zhang et al., 2023b), and the correlation ratio (Kotelnikov et al., 2023) are applied for different types of two-way marginals (categorical, continuous, and mixed). In addition to the lack of universality, this approach also suffers from the problem that correlation scores capture only limited information about the data distribution. Two attributes may have the same correlation score both in the real data and in the synthetic data, yet their underlying distributions diverge significantly—a phenomenon known as the scale invariance of correlation statistics (Wikipedia contributors, 2024).
@@
Likelihood Fitness. Some studies (Xu et al., 2019) assume the input data are generated from some known probabilistic models (e.g., Bayesian networks), thus the likelihood of synthetic data can be derived by fitting them to the priors. While likelihood fitness can naturally reflect the closeness of synthetic data to the assumed prior distribution, it is only feasible for data whose priors are known, which is inaccessible for most real-world complex datasets.
@@
Other Approaches. $\alpha$-Precision (Alaa et al., 2022) defines fidelity as the proportion that the synthetic samples are covered by real data. A synthetic record is covered if its distance to the nearest real point is within a predefined threshold. Probabilistic mean squared error (pMSE) (Snoke et al., 2018) employs a logistic regression discriminator to distinguish between synthetic and real data, using relative prediction confidence as the fidelity metric. The effectiveness of these heuristic metrics highly relies on the choice of auxiliary evaluators (e.g., threshold and discriminator), which requires careful calibration to ensure meaningful comparisons across different datasets and synthesizers.
3.2. Proposed Metric: Wasserstein Distance
@@
We opt for Wasserstein distance to measure the distribution discrepancies between synthetic data and real data. Originating from optimal transport theory (Peyré et al., 2019), the Wasserstein distance provides a structure-aware measure of the minimal amount of work required to transform one distribution into another. Formally, Let $\mathbf{P}=(p_{1},p_{2},\ldots p_{n})$ and $\mathbf{Q}=(q_{1},q_{2},\ldots q_{n})$ be the two probability distributions, and $\mathbf{C}$ be a matrix of size $n\times n$ in which $\mathbf{C}_{ij}\geq 0$ is the cost of moving an element $i$ of $\mathbf{P}$ to the element $j$ of $\mathbf{Q}$ ($\mathbf{C}_{ii}=0$ for all element $i$). The optimal transport plan $\mathbf{A}$ is defined as below:
@@
(1) $\begin{split}&\min_{\mathbf{A}}\quad\langle\mathbf{C},\mathbf{A}\rangle\\ \text{ s.t. }&\mathbf{A}\mathds{1}=\mathbf{P},\quad\mathbf{A}^{\top}\mathds{1}=\mathbf{Q},\end{split}$
@@
(2) $\mathcal{W}(\mathbf{P},\mathbf{Q})=\langle\mathbf{C},\mathbf{A}^{*}\rangle.$
@@
Definition 0.
(Wasserstein-based Fidelity Metric) Let $\mathbf{P}$ and $\mathbf{Q}$ be the marginal distributions derived from synthetic data $\mathcal{S}$ and real data $\mathcal{D}$, the fidelity of synthesis algorithm $\mathsf{A}$ is:
(3) $\mathrm{Fidelity}(\mathsf{A})\coloneqq\underset{\begin{subarray}{c}\mathbf{P}\sim\mathcal{S}\\ \mathbf{Q}\sim\mathcal{D}\end{subarray}}{\mathbb{E}}\big{[}\mathcal{W}(\mathbf{P},\mathbf{Q})\big{]},$
@@
Determining Cost Matrix. The Wasserstein distance requires the predefined cost matrix $\mathbf{C}$, which encapsulates the “cost” of transitioning from one distribution element to another. For $k$-way marginal distributions $\mathbf{P}$ and $\mathbf{Q}$, the cost matrix is formulated by summing the pairwise distances between corresponding elements:
(4) $\mathbf{C}_{ij}=\sum_{r}d(v_{i}^{r},v_{j}^{r}).$
Here, $v_{i},v_{j}\in\mathbb{R}^{k}$ are the element located in $i$ and $j$ in $k$-way probability distributions. The distance $d(\cdot,\cdot)$ is tailored to the nature of the attributes, differing for numerical and categorical values:
(5) $d(v_{i}^{r},v_{j}^{r})=\begin{cases}||v_{i}^{r}-v_{j}^{r}||_{1}&\text{if numerical}\\ \infty,&\forall v_{i}^{r}\neq v_{j}^{r},\text{if categorical}.\end{cases}$
This means we use $l_{1}$ distance for numerical values and consider the strict match for categorical values (i.e., cost for any different categorical values is infinity). Note that it is also feasible to assign semantic distance for categorical attributes (Li et al., 2021), we omit it because it requires specific context for optimization and most synthesizers do not model the semantics in tabular data.
Implementations. In practice, the real dataset $\mathcal{D}$ can be designated as either $\mathcal{D}_{train}$ or $\mathcal{D}_{test}$, depending on the specific requirements of analysis. The computation of Wasserstein distance involves solving the optimization problem in Equation 1, which is an instance of linear programming problems (LP) (Peyré et al., 2019). There are many open-source libraries like CVXPY (Diamond and Boyd, 2016) and OPT (Flamary et al., 2021) that can be used to solve LP reasonably fast [^4] If the computation is still too slow, Sinkhorn distance (Cuturi, 2013), Sliced-Wasserstein distance (Bonneel et al., 2015) and random sampling can be used to approximate the Wasserstein distance. We implement these methods in our framework and detailed in technical report version.. In our evaluation, we use OPT package (Flamary et al., 2021) to compute all the one-way and two-way marginals and use the mean as the final fidelity score.
Merits of Wasserstein-based Fidelity Metric. Wasserstein distance offers several distinct advantages for fidelity evaluations: (i) Faithfulness. It serves as a natural and structure-aware statistic measurement for analyzing distribution discrepancies, which generalizes existing metrics such as TVD and contingency similarity (Patki et al., 2016; Zhang et al., 2023b). (ii) Universality. It can accommodate both numerical and categorical attributes and extend to any multivariate marginals under the same criterion, facilitating the evaluation of heterogeneous types of marginals.
4. Privacy Evaluation
@@
4.1. Existing Metrics and Limitations
@@
Similarity-based Privacy Metrics. Similarity-based approaches measure the empirical privacy risk of synthetic data by comparing the closeness with real data, which have become the conventional privacy evaluation metrics for HP synthesizers (Lu et al., 2019; Ganev and De Cristofaro, 2023; Jordon et al., 2021). Among all heuristics, the Distance to Closest Records (DCR) is the most widely used both in academia (Zhao et al., 2021; Yale et al., 2019; Kotelnikov et al., 2023; Zhang et al., 2023b) and industry (AWS, 2022; Gretel, 2023; Mendelevitch and Lesh, 2021). DCR looks at the distribution of the distances from each synthetic data point to its nearest real one, and uses the 5th percentile (or the mean) of this distribution as the privacy score. A small score indicates that the synthetic dataset is too close (similar) to real data, signaling a high risk of information leakage.
We point out two drawbacks of the DCR metrics. First, DCR measures privacy loss averaged over the data points, whereas it is crucial to protect the privacy of every individual. When measuring the privacy leakage across different individuals, one needs to ensure that the worst-case leakage is bounded. It is unacceptable to use a mechanism that sacrifices the privacy of some individuals even though the protection averaged over the population is good. This point is illustrated by the fact that the re-identification of one or a few individuals is commonly accepted as privacy breaches (Li et al., 2013; Dwork, 2006; UN, 2023; OECD, 2023). DCR uses the 5th percentile (or mean) proximity to real data as a measurement, and fails to provide worst-case protection.
Second, DCR overestimates the privacy risks when data points are naturally clustered close together. As illustrated by discussions about differential privacy (Dwork et al., 2014; Li et al., 2017), leaking information regarding an individual should not be considered a privacy violation if the leakage can occur even if the individual’s data is not used. Analogously, having some synthetic data very close to real ones does not mean worse privacy if this situation can occur even if each data point is removed. Consider, for example, a dataset that is a mixture of two Gaussians with small standard deviations. A good synthetic dataset is likely to follow the same distribution, and has many data points very close to the real ones. DCR interprets this closeness as a high privacy risk, overlooking the fact that the influence of any individual training instance on synthetic data is insignificant.
@@
Another concern of the similarity-based privacy metrics is the risk of additional information leakage when the adversary gets access to these metric values. Since these metrics directly measure the distance between synthetic data points and real ones, these metric values provide direct information about the real data points. Recent research (Ganev and De Cristofaro, 2023) has demonstrated the feasibility of reconstructing significant portions of training data when given black-box access to a synthesizer and an oracle for computing the DCR. This revelation is a stark warning to practitioners relying on such metrics.
@@
Membership Inference Attacks. Another way to empirically verify the privacy of machine learning models is membership inference (MI) attacks (Shokri et al., 2017), which have been extensively studied on discriminative models. Some studies (Stadler et al., 2022) try to perform the attack on data synthesis by utilizing handcrafted features extracted from synthetic data distribution to train shadow models. Another work (Hyeong et al., 2022) approximates the likelihood of target records with density estimation and uses it as the confidence score for membership inference. However, as pointed out in (Giomi et al., 2023) and (Qian et al., 2023), current MI attacks on synthetic data either depend on unrealistic settings or are too weak to differentiate the membership disclosure risks among various synthesizers (Dionysiou and Athanasopoulos, 2023). Therefore, existing synthesis algorithms rarely use MI attacks for privacy evaluation (Qian et al., 2023).
4.2. Proposed Metric: Membership Disclosure Score (MDS)
We propose a new privacy metric, termed membership disclosure score (MDS), to estimate the membership disclosure risks of data synthesizers. MDS follows the same intuition behind DCR, namely, including a data point $x$ in the data synthesizing process may result in the generation of data points that are close to $x$, and the degree of closeness is related to the degree of privacy leakage. However, we avoid the two drawbacks of DCR identified above. First, given an input dataset $\mathcal{D}$ and synthesizer $\mathsf{A}$, we quantify the disclosure risk of one record $x\in\mathcal{D}$ as follows:
@@ 4.2. Proposed Metric: Membership Disclosure Score (MDS)
(6) $\mathrm{DS}(x,\mathsf{A},\mathcal{D})\coloneqq\underset{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime}\subset\mathcal{D}\\ {\mathcal{H}\cap\mathcal{H}^{\prime}=\{x\}}\end{subarray}}{\mathbb{E}}\big{[}||\text{NN}(\mathcal{S},x)-\text{NN}(\mathcal{S}^{\prime},x)||_{1}\big{]},$
@@
where the $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are subsets of training samples sampled from the dataset $\mathcal{D}$, and they only differ in $x$. $\mathcal{S}$ and $\mathcal{S}^{\prime}$ represent the synthetic datasets when the synthesizer $\mathsf{A}$ is trained on $\mathcal{H}$ and $\mathcal{H}^{\prime}$. The term $\text{NN}(\mathcal{S},x)$ denotes the nearest neighbor (under $l_{1}$ distance) for record $x$ within the synthetic dataset $\mathcal{S}$. We also use $l_{1}$ distance (empirically we find that the difference between using $l_{1}$ and $l_{2}$ distance is negligible) to measure the discrepancy resulting from training the model with or without the record $x$.
$\mathrm{DS}(x,\mathsf{A},\mathcal{D})$ addresses DCR’s second drawback identified above (over-estimating leakage) by quantifying how much including $x$ changes the distance between $x$ and the closest synthetic data. If including $x$ results in records much closer to $x$ to be generated, then the disclosure risk is high. If records close to $x$ are generated whether or not $x$ is included, then the disclosure risk for $x$ is low.
We then define the membership disclosure risk of synthesizer $\mathsf{A}$ on dataset $\mathcal{D}$ to be the maximum disclosure risk among all records. This addresses DCR’s first drawback identified above.
Definition 0.
@@
(Membership Disclosure Score) Let $\mathrm{DS}(x,\mathsf{A},\mathcal{D})$ be the disclosure risk of record $x$ in real dataset $\mathcal{D}$ for synthesis algorithm $\mathsf{A}$. The membership disclosure score of $\mathsf{A}$ is given by:
(7) $\mathrm{MDS}(\mathsf{A})\coloneqq\max_{x\in\mathcal{D}}~{}\mathrm{DS}(x,\mathsf{A},\mathcal{D}).$
@@
Here we omit the notion of dataset $\mathcal{D}$ for simplicity. This metric signifies the maximum disclosure risk across the entire training dataset, which indicates that MDS is a worst-case privacy measurement. Intuitively, for a privacy-preserving synthesizer, the change of output (synthetic data) should remain insignificant by any individual training samples, resulting in a low MDS. It is worth mentioning that the disclosure score is calculated based on the difference between neighboring synthetic data, which ensures that real data information remains undisclosed (unlike DCR).
@@
Implementation. Directly computing this score entails retraining many models for each record, which is computationally prohibitive. Alternatively, we train $m$ models once on random subsets of $\mathcal{D}$, and compute the disclosure score for all instances using the same set of $m$ synthesis models (statistically each record is trained on half of the models, while the other half are not). This approach provides an efficient and practical approach to estimating MDS in real-world scenarios. Empirically we find that $m=80$ is sufficient to accurately estimate MDS for all synthesizers.
Limitations of MDS. In Section 7.5, we show that MDS can differentiate the privacy risks of different synthesizers where DCR cannot. However, MDS has its own limitations and is not designed to replace metrics based on differential privacy or MI attacks. Specifically, MDS applies only to algorithms that learn the distribution of the input records and then synthesize from the distribution. There exist pathological synthesizers for which MDS is inappropriate. One example of such a synthesizer is the one that simply adds the same large constant to each record in the input dataset. In datasets where the records are far apart, this results in a large MDS (meaning low disclosure risks) yet completely reveals the dataset, assuming that the mechanism of the synthesizer is known. However, most (if not all) practical synthesizers are not such pathological ones, i.e., they learn the distribution and then generate records from it. In summary, whenever DCR is used as a heuristic privacy metric, we believe that it is better to use MDS instead. More discussion about MDS is included in our technical report version.
5. Utility Evaluation